OpenStudy (666potatoes):

Rationalize denominator: 4/3 + 5i Actual steps please. <3

OpenStudy (666potatoes):

I know that you have to multiply the top and bottom by 3+5i and I got 12+20i/16+30i but I have idea where to go from there.

OpenStudy (anonymous):

multiply by the conjugate

OpenStudy (tkhunny):

First, you must write what you mean. As it is written, you have \(\dfrac{4}{3} + 5i\) and the denominator is already rational. \(\dfrac{4}{3+5i}\cdot\dfrac{3-5i}{3-5i} = \dfrac{12 - 20i}{3^{2} + 5^{2}}\). Remember that \(i^{2} = -1\)

OpenStudy (666potatoes):

It's actually 4/(3+5i)